# ChronoMagic-Bench: A Benchmark for Metamorphic Evaluation of Text-to-Time-lapse Video Generation

### v1

@@ Abstract
We propose a novel text-to-video (T2V) generation benchmark, ChronoMagic-Bench111 ”Chrono” is derived from the Greek word ”chronos”, which means ”time”., to evaluate the temporal and metamorphic capabilities of the T2V models (e.g. Sora [8] and Lumiere [3]) in time-lapse video generation. In contrast to existing benchmarks that focus on visual quality and textual relevance of generated videos, ChronoMagic-Bench focuses on the models’ ability to generate time-lapse videos with significant metamorphic amplitude and temporal coherence. The benchmark probes T2V models for their physics, biology, and chemistry capabilities, in a free-form text query. For these purposes, ChronoMagic-Bench introduces 1,649 prompts and real-world videos as references, categorized into four major types of time-lapse videos: biological, human-created, meteorological, and physical phenomena, which are further divided into 75 subcategories. This categorization ensures a comprehensive evaluation of the models’ capacity to handle diverse and complex transformations. To accurately align human preference with the benchmark, we introduce two new automatic metrics, MTScore and CHScore, to evaluate the videos’ metamorphic attributes and temporal coherence. MTScore measures the metamorphic amplitude, reflecting the degree of change over time, while CHScore assesses the temporal coherence, ensuring the generated videos maintain logical progression and continuity. Based on the ChronoMagic-Bench, we conduct comprehensive manual evaluations of ten representative T2V models, revealing their strengths and weaknesses across different categories of prompts, and providing a thorough evaluation framework that addresses current gaps in video generation research. Moreover, we create a large-scale ChronoMagic-Pro dataset, containing 460k high-quality pairs of 720p time-lapse videos and detailed captions. Each caption ensures high physical pertinence and large metamorphic amplitude, which have a far-reaching impact on the T2V generation community.
@@ 1 Introduction
Text-to-video (T2V) generative models [84, 86, 40, 22, 69, 62, 28, 79] have been developed rapidly recently. As the number of models continues to grow, there is an urgent need for evaluation methods that align with human perception, accurately reflecting the specific strengths and weaknesses of each model, thereby enabling community to adopt architectures that meet their requirements more easily.
However, the current T2V benchmarks [43, 61, 59, 75, 26, 28] primarily assess the capability of generating general videos instead of time-lapse videos, failing to reflect the extent of physical priors encoded by the models. Additionally, the evaluation metrics they use mainly focus on visual quality and textual relevance, from early metrics including FID [24], FVD [64], and CLIPScore [23] to more recent ones such as UMTScore [43], T2VQA [30], and UMT-FVD [43], almost all of which overlook two other crucial aspects of videos: metamorphic amplitude and temporal coherence. These limitations hinder the development of T2V models in generating videos with rich physical content.
@@
| ChronoMagic-Bench | Time-lapse | ✓ | ✓ | ✓ | ✓ |
Due to the greater metamorphic amplitude and temporal coherence of time-lapse videos, they contain more physical priors compared to general videos [79]. Therefore, to address the aforementioned issues, we introduce a benchmark called ChronoMagic-Bench for Metamorphic Evaluation of Time-Lapse Text-to-Video Generation, which provides a comprehensive evaluation system for T2V. We specifically design four major categories for time-lapse videos, including biological, human-created, meteorological, and physical videos, and extend these to 75 subcategories. Based on this, we construct ChronoMagic-Bench, comprising 1,649 prompts and their corresponding reference time-lapse videos. As shown in Table 1, In contrast to existing benchmarks [59, 75, 43, 26, 72, 61], ChronoMagic-Bench emphasizes generating videos with high persistence and strong variation, i.e., metamorphic videos with high physical prior content. Additionally, we develop MTScore for evaluating metamorphic amplitude and CHScore for temporal coherence to address the deficiencies in evaluation metrics and perspectives. With ChronoMagic-Bench, we conduct comprehensive qualitative and quantitative evaluations of almost all open-source T2V models, enabling analysis of their strengths and weaknesses. The results highlight the weaknesses of existing T2V models, including 1) failure by almost all models to generate time-lapse videos with large variations; 2) poor adherence to prompts (thus necessitating multiple inferences to achieve satisfactory results); and 3) flickering even though the visual quality of single frames may be high (indicating poor temporal coherence).
Furthermore, we have meticulously curated the dataset ChronoMagic-Pro to provide the community with the first large-scale T2V dataset specifically designed for time-lapse video generation with higher physical prior content. ChronoMagic-Pro stands out from previous T2V datasets [75, 2, 70, 15, 68] as it comprises time-lapse videos (e.g., ice melting and flowers blooming) characterized by strong physical characteristics, high persistence, and variability. Considering the domain differences between time-lapse videos and general videos, we propose an automatic time-lapse video collection framework to ensure the integrity of video content and improve annotation quality.
In summary, the contributions of this work are as follows:
i) A New T2V Benchmark. We introduce ChronoMagic-Bench to evaluate the existing T2V models from visual quality, textual relevance, metamorphic amplitude, and temporal coherence.
@@
iii) New Insights for T2V Model Selection. Our evaluations using ChronoMagic-Bench provide crucial insights into the strengths and weaknesses of various open/close-source T2V models.
iv) A Large-Scale Time-lapse Video-Text Dataset. We create ChronoMagic-Pro, a dataset with 460k high-quality 720p time-lapse videos and detailed captions, promoting advances in T2V research.
@@ 2 Related Work
Automatic Metrics for Text-to-Video Generation. Existing benchmarks [26, 31, 73, 34, 54] typically utilize Frechet Inception Distance (FID) [24], Frechet Video Distance (FVD) [64], CLIPScore [23], or their improved versions to assess the visual quality and text relevance of generated videos. For example, FETV [43] enhances FVD and CLIPScore within the UMT [36] feature space, resulting in UMT-FVD and UMTScore. Additionally, the CLIPScore feature extractor can be replaced with BLIP [82] to evaluate the relevance between text and generated content. To the best of our knowledge, existing T2V benchmarks [42, 43, 78, 18, 48] mainly assess these two aspects, with prompts based on general videos. This means that temporal coherence and metamorphic amplitude in videos have been overlooked, leading to the absence of automated metrics that indirectly reflect the physical content encoded by video models. Although some work [43, 26, 42] assess coherence, they are based on feature space or human evaluation, which is expensive and not sufficiently intuitive. Therefore, we propose the Metamorphic Score (MTScore) and Coherence Score (CHScore) to measure the metamorphic degree and temporal coherence of videos, filling this gap in the field.
@@ 3.1 Benchmark Construction
Prompt Construction. To comprehensively test the time-lapse video generation capabilities of existing T2V models, the designed text prompts need to cover as many metamorphic types as possible, and the corresponding reference videos must be of relatively high quality. Manual construction is impractical; therefore, to build a T2V benchmark rich in visual concepts, we first manually created a search term database suitable for diverse and broadly applicable time-lapse videos. We then counted the number of videos obtainable for each search term and filtered them based on frequency, resulting in a search database containing 75 categories of time-lapse videos. Additionally, since there are four major nature phenomena: biological covers all content related to living organisms, such as plant growth, animal activities, microbial movement, etc. Human-created includes all objects created or influenced by human activities, such as the construction process of buildings, urban traffic flow, etc. Meteorological includes all content related to meteorological phenomena, such as cloud movement, storm formation, etc. Physical includes all content related to non-biological physical phenomena, such as water flow, volcanic eruptions, etc. We divide the 75 subcategories into four major categories (biological, human-created, meteorological, and physical), as shown in Figure 1. Then, we use the search terms to crawl 20 high-quality videos for each category from video platforms. Finally, we used GPT-4o [1] to accurately caption these 1,649 videos and treated these captions as the text prompts for the benchmark. For more details about benchmark construction, please refer to Appendix E.
@@
Benchmark Statistics. We collect a total of 1,649 prompts with corresponding videos and categories, the specific data distribution is shown in Figure 2, indicating that 75 categories have a comparable number of test cases to reflect the time-lapse video generation capabilities of different models accurately. Each data sample in ChronoMagic-Bench consists of four elements: prompt , reference video , sub-category , and major category . Since existing T2V models typically use CLIP as the text encoder, which supports a maximum input of 77 tokens, we have limited the length of to within 77 tokens for general applicability, as shown in Figure 3(a). Although the length is limited, the diversity remains rich. By comparing the main words in the word cloud, as shown in Figure 3(b), it is observed that terms related to time-lapse videos such as "transitioning," "progressing," "increasing," and "gradually" appear most frequently. These terms significantly highlight ChronoMagic-Bench’s focus on large metamorphic amplitude, strong persistence of changes, and high physical content. In addition, words from four major categories are distributed, such as biological (seed, butterfly, etc.), human-created (Minecraft, traffic, etc), meteorological (sunset, tide, etc), and physical (burning, explosion, etc). For detailed explanations of the 75 subcategories, please refer to the Appendix E.
@@ 3.2 New Automatic Metrics
Due to the strong instruction-following capability and world-understanding ability of GPT-4o, it can partially replace humans. For the fine-grained score (GPT4o-MTScore), we use GPT-4o as the evaluator. Specifically, we set a 5-point evaluation standard, then uniformly sample frames and input them into GPT-4o[1] to obtain the score. More details are provided in Appendix B.1.
Temporal Coherence Score. Temporal coherence is crucial for time-lapse videos because they span a large time range. Current benchmarks assess coherence either through questionnaires [43] or by employing methods based on feature space calculations [26, 42]. The former approach is time-consuming, whereas the latter lacks intuitiveness and does not support visualization. Therefore, we developed the Coherence Score (CHScore) based on a video tracking model [27]. More details are provided in the Appendix B.2. Specifically, we first process input video using the pre-trained model with grid size and threshold to get . Then, we count the number of missing tracking points in each frame, and the change in missed points between consecutive frames :
where , represents the position of the frame, and identifies different tracking points. Based on these, we then calculate the , which represents the average proportion of missed points per frame in the video. And the , which measures the variation in the number of missed points between consecutive frames, indicating frame-to-frame coherence:
where , is the mean of , is the total number of frames, is the number of points per frame, and indicates the visibility of point in frame . In addition, we need to calculate the , which indicates the ratio of frames that need to be cut to the total number of frames, reflecting the extent of video editing required. And the , which indicates the number of consecutive changes in missed points exceeding the threshold, indicating frequent large-scale instability in point tracking:
@@ 3.3 Application Scope
| WebVid-10M [2] | General | 10M | 360p | Video-Text | 18.7s | 52K |
| InternVid [70] | General | 234M | 720p | Video-Text | 11.9s | 760K |
| Panda-70M [15] | General | 70M | 720p | Video-Text | 8.5s | 166K |
| HD-VG-130M [68] | General | 130M | 720p | Video-Text | 4.9s | 178K |
@@
| ChronoMagic-Pro | Time-lapse | 460K | 720p | Video-Text | 234.1s | 30K |
| ChronoMagic-ProH | Time-lapse | 150K | 720p | Video-Text | 190.2s | 8K |
@@ 4 ChronoMagic-Pro
Multi-Aspect Data Curation. As previously mentioned, existing large-scale text-video datasets primarily consist of general videos with limited physical information content, restricting open-source models [5, 59, 22] to generating only general videos rather than time-lapse videos. To address this, we construct the first large-scale time-lapse video dataset by collecting time-lapse videos based on the search terms outlined in Section 3.1, ultimately obtaining 66,226 original videos. Following the Panda70m method [15], we split these videos to produce 460K semantically consistent single-scene video clips (ChronoMagic-Pro). Finally, we utilize the video annotation strategy similar to MagicTime [79], replacing GPT-4V [1] with the open-source ShareGPT4Video [12] to reduce computational overhead while ensuring high-quality video captions. Additionally, we employ a video retrieval model [71] to filter out low-quality videos, resulting in 150K video clips of higher purity and quality (ChronoMagic-ProH). Verification experiment are shown in the Appendix D.3.
Dataset Statistics. We collect time-lapse videos from 75 categories manually set by the human, with proportions being roughly similar. Some samples can be found in Appendix C.3. Unlike before, ChronoMagic-Pro is the first high-quality large-scale time-lapse T2V dataset, which contains more physical knowledge than general videos, as shown in Table 2. As shown in Figure 4, in terms of duration, more than half (53.3%) of the videos have a duration of 0-15 seconds, a quarter (27.1%) are longer than 60 seconds, 12.1% are between 15-30 seconds, and the remaining videos are distributed between 30-60 seconds. Regarding resolution, 97% are high resolution (720P), 2% are ultra-high resolution (1080P), and the remaining videos have lower resolutions ranging from 360P to 480P. As the number of words accepted by the text encoder increases, we require the generated captions to be as detailed as possible, with 95% of captions containing more than 100 words. For aesthetic score [57], 73% videos receive high scores ranging from 4 to 6. 14% of the videos had aesthetic indicators exceeding 6, and only a small portion of the videos score below 3. This indicates that the quality of most videos is high. For the word distribution of the generated captions, please refer to Appendix C.2. Similar to Figure 3, ChronoMagic-Pro mainly focuses on changes (gradually, progressing, increasing, etc.), processes spanning a large amount of time, such as flower blooming and ice melting. In addition, the categories, durations, resolutions, and word count ranges of ChronoMagic-ProH are similar to those of ChronoMagic-Pro; however, ChronoMagic-ProH exhibits superior quality and purity.
@@
| ModelScopeT2V [66] | Arxiv’23 | U-Net | 194.77 | 2.909 | 0.401 | 11.03 | 2.86 |
| ZeroScope [62] | CVPR’23 | U-Net | 227.02 | 2.350 | 0.400 | 25.13 | 2.09 |
| T2V-zero [28] | ICCV’23 | U-Net | 209.66 | 2.661 | 0.400 | 1.68 | 2.55 |
| LaVie [69] | Arxiv’23 | U-Net | 166.97 | 2.763 | 0.346 | 8.60 | 2.46 |
| AnimateDiff [22] | ICLR’24 | U-Net | 197.89 | 2.944 | 0.467 | 11.36 | 2.62 |
| VideoCrafter2 [10] | Arxiv’24 | U-Net | 178.45 | 2.753 | 0.433 | 8.27 | 2.68 |
| MCM [80] | Arxiv’24 | U-Net | 202.08 | 2.33 | 0.417 | 14.08 | 3.04 |
| MagicTime [79] | Arxiv’24 | U-Net | 257.56 | 1.916 | 0.478 | 10.66 | 3.13 |
| Latte [45] | Arxiv’24 | DiT | 192.12 | 2.111 | 0.363 | 13.81 | 2.20 |
| OpenSora 1.1 [86] | Github’24 | DiT | 195.43 | 2.678 | 0.444 | 10.03 | 2.52 |
| OpenSora 1.2 [86] | Github’24 | DiT | 166.92 | 2.781 | 0.375 | 4.69 | 2.56 |
| OpenSoraPlan v1.1 [40] | Github’24 | DiT | 188.53 | 2.421 | 0.327 | 10.35 | 2.19 |

### v2

@@ Abstract
We propose a novel text-to-video (T2V) generation benchmark, ChronoMagic-Bench111”Chrono” derives from the Greek word ”chronos”, which means ”time”., to evaluate the temporal and metamorphic knowledge skills in time-lapse video generation of the T2V models (e.g. Sora [9] and Lumiere [4]). Compared to existing benchmarks that focus on visual quality and text relevance of generated videos, ChronoMagic-Bench focuses on the models’ ability to generate time-lapse videos with significant metamorphic amplitude and temporal coherence. The benchmark probes T2V models for their physics, biology, and chemistry capabilities, in a free-form text control. For these purposes, ChronoMagic-Bench introduces 1,649 prompts and real-world videos as references, categorized into four major types of time-lapse videos: biological, human creation, meteorological, and physical phenomena, which are further divided into 75 subcategories. This categorization ensures a comprehensive evaluation of the models’ capacity to handle diverse and complex transformations. To accurately align human preference on the benchmark, we introduce two new automatic metrics, MTScore and CHScore, to evaluate the videos’ metamorphic attributes and temporal coherence. MTScore measures the metamorphic amplitude, reflecting the degree of change over time, while CHScore assesses the temporal coherence, ensuring the generated videos maintain logical progression and continuity. Based on the ChronoMagic-Bench, we conduct comprehensive manual evaluations of ten representative T2V models, revealing their strengths and weaknesses across different categories of prompts, providing a thorough evaluation framework that addresses current gaps in video generation research. More encouragingly, we create a large-scale ChronoMagic-Pro dataset, containing 460k high-quality pairs of 720p time-lapse videos and detailed captions. Each caption ensures high physical content and large metamorphic amplitude, which have a far-reaching impact on the video generation community. 222The source data and code are publicly available on https://pku-yuangroup.github.io/ChronoMagic-Bench.
@@ 1 Introduction
Text-to-video (T2V) generative models [88, 90, 41, 22, 71, 64, 28, 83] have developed rapidly recently. As the number of models continues to grow, there is an urgent need for evaluation methods that align with human perception, accurately reflecting the specific strengths and weaknesses of each model, thereby enabling the community to more easily select architectures that meet their requirements.
However, the current T2V benchmarks [45, 63, 61, 78, 26, 28] primarily assess the capability of generating general videos instead of time-lapse videos, failing to reflect the extent of physical priors encoded by the models. Additionally, the evaluation metrics they use mainly focus on visual quality and textual relevance, from early metrics like FID [24], FVD [66], and CLIPScore [23] to more recent ones like UMTScore [45], T2VQA [31], and UMT-FVD [45], all of which overlook two other crucial aspects of videos: metamorphic amplitude and temporal coherence. These limitations hinder the development of T2V models in generating videos with rich physical content.
@@
| ChronoMagic-Bench (Ours) | Time-lapse | ✓ | ✓ | ✓ | ✓ |
Due to the greater metamorphic amplitude and temporal coherence of time-lapse videos, they contain more physical priors compared to general videos [83]. Therefore, to address the aforementioned issues, we introduce a benchmark called ChronoMagic-Bench for Metamorphic Evaluation of Time-Lapse Text-to-Video Generation, which provides a comprehensive evaluation system for T2V. We specifically designed four major categories for time-lapse videos, including biological, human creation, meteorological, and physical, and extended these to 75 subcategories. Based on this, we constructed ChronoMagic-Bench, comprising 1,649 prompts and their corresponding reference time-lapse videos. As shown in Table 1, compared to existing benchmarks [61, 78, 45, 26, 74, 63], ChronoMagic-Bench emphasizes generating videos with high persistence and strong variation, i.e., metamorphic videos with high physical prior content. Additionally, we developed MTScore for evaluating metamorphic amplitude and CHScore for temporal coherence to address the deficiencies in evaluation metrics and perspectives. With ChronoMagic-Bench, we conducted comprehensive qualitative and quantitative evaluations of almost all open&closed-source T2V models, enabling analysis of their strengths and weaknesses. The results highlighted the weaknesses of these models, including (1) almost all models fail to generate time-lapse videos with large variations; (2) poor adherence to prompts, necessitating multiple inferences to achieve satisfactory results; (3) the visual quality of single frame may be high, but flickering may occur, indicating poor temporal coherence.
Furthermore, we have meticulously curated the dataset ChronoMagic-Pro to provide the community with the first large-scale T2V dataset specifically designed for time-lapse video generation with higher physical prior content. ChronoMagic-Pro stands out from previous T2V datasets [78, 2, 72, 16, 70] as it comprises time-lapse videos (e.g., ice melting and flowers blooming) characterized by strong physical characteristics, high persistence, and variability. Considering the domain differences between time-lapse videos and general videos, we proposed an automatic time-lapse video collection framework to maintain video purity and improve annotation quality.
The contributions of this work are as follows:
i) New T2V Benchmark. We introduce ChronoMagic-Bench for comprehensive evaluation of T2V models, focusing on visual quality, text relevance, metamorphic amplitude, and temporal coherence.
@@
iii) New Insights for T2V Model Selection. Our evaluations using ChronoMagic-Bench provide crucial insights into the strengths and weaknesses of various T2V models.
iv) Large-Scale Time-lapse Video-Text Dataset. We create ChronoMagic-Pro, a dataset with 460k high-quality 720p time-lapse videos and detailed captions, promoting advances in T2V research.
@@ 2 Related Work
Automatic Metrics for Text-to-Video Generation. Existing benchmarks [26, 32, 75, 35, 56] typically utilize Frechet Inception Distance (FID) [24], Frechet Video Distance (FVD) [66], CLIPScore [23], or their improved versions to assess the visual quality and text relevance of generated videos. For example, FETV [45] enhances FVD and CLIPScore within the UMT [37] feature space, resulting in UMT-FVD and UMTScore. Additionally, the CLIPScore feature extractor can be replaced with BLIP [86] to evaluate the relevance between text and generated content. To the best of our knowledge, existing T2V benchmarks [44, 45, 82, 18, 50] mainly assess these two aspects, with prompts based on general videos. This means that temporal coherence and metamorphic amplitude in videos have been overlooked, leading to the absence of automated metrics that indirectly reflect the physical content encoded by video models. Although [45, 26, 44] assess coherence, they are based on feature space or human evaluation, which is expensive and not sufficiently intuitive. Therefore, we propose the Metamorphic Score (MTScore) and Coherence Score (CHScore) to measure the metamorphic degree and temporal coherence of videos, filling this gap in the field.
@@ 3.1 Benchmark Construction
Prompt Construction. To comprehensively evaluate the time-lapse video generation capabilities of existing T2V models, the designed text prompts need to cover as many metamorphic types as possible, and the corresponding reference videos must be of relatively high quality. Manual construction is impractical; therefore, to build a T2V benchmark rich in visual concepts, we first manually created a search term database suitable for diverse and broadly applicable time-lapse videos. We then counted the number of videos obtainable for each search term and filtered them based on frequency, resulting in a search database containing 75 categories of time-lapse videos. Additionally, since there are four major nature phenomena: biological covers all content related to living organisms, such as plant growth, animal activities, microbial movement, etc. Human creation includes all objects created or influenced by human activities, such as the construction process of buildings, urban traffic flow, etc. Meteorological includes all content related to meteorological phenomena, such as cloud movement, storm formation, etc. Physical includes all content related to non-biological physical phenomena, such as water flow, volcanic eruptions, etc. We divide the 75 subcategories into four major categories (biological, human creation, meteorological, and physical), as shown in Figure 1. We then utilized a search engine to crawl 20 or more high-quality videos from video platforms for each category, ultimately gathering a total of 1,649 videos. Finally, we use GPT-4o [1] to accurately caption these videos and treat these captions as the text prompts for the benchmark. For more details about benchmark construction, please refer to Appendix E.
@@
Benchmark Statistics. We collect a total of 1,649 prompts with corresponding videos and categories, the specific data distribution is shown in Figure 2, indicating that 75 categories have a comparable number of test cases to reflect the time-lapse video generation capabilities of different models accurately. Each data sample in ChronoMagic-Bench consists of four elements: prompt , reference video , sub-category , and major category . Since existing T2V models typically use CLIP as the text encoder, which supports a maximum input of 77 tokens, we have limited the length of to within 77 tokens for general applicability, as shown in Figure 3(a). Although the length is limited, the diversity remains rich. By comparing the main words in the word cloud, as shown in Figure 3(b), it is observed that terms related to time-lapse videos such as "transitioning," "progressing," "increasing," and "gradually" appear most frequently. These terms significantly highlight ChronoMagic-Bench’s focus on large metamorphic amplitude, strong persistence of changes, and high physical content. In addition, words from four major categories are distributed, such as biological (seed, butterfly, etc.), human creation (Minecraft, traffic, etc), meteorological (sunset, tide, etc), and physical (burning, explosion, etc). For detailed explanations of the 75 subcategories, please refer to the Appendix E.
@@ 3.2 New Automatic Metrics
Due to the strong instruction-following capability and world-understanding ability of GPT-4o, it can partially replace humans. For the fine-grained score (GPT4o-MTScore), we use GPT-4o as the evaluator. Specifically, we set a 5-point evaluation standard, then uniformly sample frames and input them into GPT-4o[1] to get the score. More implementation details are provided in Appendix B.
Temporal Coherence Score. Temporal coherence is crucial for time-lapse videos because they span a large time range. Current benchmarks assess coherence either through questionnaires [45] or by employing methods based on feature space calculations [26, 44]. The former approach is time-consuming, whereas the latter lacks intuitiveness and does not support visualization. Therefore, we developed the Coherence Score (CHScore) based on a video tracking model [27] as shown in Algorithm 1. More details are provided in the Appendix B.2. First, we process the input video using a pre-trained model with grid size and threshold to obtain (the visibility of point in frame ). Next, we count the number of missing tracking points in each frame and the change in missed points between consecutive frames . To make the CHScore robust to the temporally coherent disappearance of points, we further calculate the direction of camera/object movement based on the tracking points across all frames. If the tracking point j of frame i disappears in the far direction, it is not included in the calculation of . We then calculate the average proportion of missed points per frame , indicating the overall visibility issue across the video. Following this, we compute the variation in the number of missed points between consecutive frames , measuring frame-to-frame coherence. We also determine the ratio of frames that need to be cut , reflecting the extent of video editing required, and count the number of consecutive changes in missed points exceeding the threshold , indicating frequent large-scale instability in point tracking. Additionally, we measure the maximum continuous change in missed points , highlighting the most severe continuity breaks in the video. Finally, we integrate these metrics to calculate the Coherence Score (CHScore). In the subsequent section, the actual CHScore is scaled by 0.1 to provide a more concise representation. Further details can be found in the appendix B.2.
@@ 3.3 Application Scope
| WebVid-10M [2] | General | 10M | 360p | Video-Text | 18.72s | 52K |
| InternVid [72] | General | 234M | 720p | Video-Text | 11.90s | 760.3K |
| Panda-70M [16] | General | 70M | 720p | Video-Text | 8.50s | 166.8K |
| HD-VG-130M [70] | General | 130M | 720p | Video-Text | 4.93s | 178K |
@@
| ChronoMagic-Pro | Time-lapse | 460K | 720p | Video-Text | 234s | 30K |
@@ 4 ChronoMagic-Pro
Multi-Aspect Data Curation. As previously mentioned, existing large-scale text-video datasets primarily consist of general videos with limited physical information content, restricting open-source models [6, 61, 22] to generating only general videos rather than time-lapse videos. To address this, we construct the first large-scale time-lapse video dataset by collecting time-lapse videos based on the search terms outlined in Section 3.1, ultimately obtaining 66,226 original videos. Following the Panda70m method [16], we split these videos to produce 460K semantically consistent single-scene video clips. Finally, we utilize the video annotation strategy similar to MagicTime [83], replacing GPT-4V [1] with the open-source ShareGPT4Video [13] to reduce computational overhead while ensuring high-quality video captions. We conduct a verification experiment on the dataset in the Appendix D.3. More details about dataset construction are provided in Appendix C.
Dataset Statistics. We collected time-lapse videos from 75 categories manually set by the human, with proportions being roughly similar. Some samples can be found in the Appendix C.4. ChronoMagic-Pro is the first high-quality large-scale time-lapse T2V dataset, which contains more physical knowledge than general videos, as shown in Table 2. As shown in Figure 4, in terms of duration, more than half (53.3%) of the videos have a duration of 0-15 seconds, a quarter (27.1%) are longer than 60 seconds, 12.1% are between 15-30 seconds, and the remaining videos are distributed between 30-60 seconds. Regarding resolution, 97% are high resolution (720P), 2% are ultra-high resolution (1080P), and the remaining videos have lower resolutions ranging from 360P to 480P. As the number of words accepted by the text encoder increases, we require the generated captions to be as detailed as possible, with 95% of captions containing more than 100 words. For aesthetic score [59], 73% videos get high scores ranging from 4 to 6. 14% of the videos had aesthetic indicators exceeding 6, and only a small portion of the videos scored below 3. This indicates that the quality of most videos is high. For the word distribution of the generated captions, please refer to Appendix C.3. Similar to Figure 3, ChronoMagic-Pro mainly focuses on changes (gradually, progressing, increasing, etc.), processes spanning a large amount of time, such as flower blooming and ice melting.
@@
| ModelScopeT2V [68] | Arxiv’23 | U-Net | 194.77 | 2.909 | 0.401 | 61.07 | 2.86 |
| ZeroScope [64] | CVPR’23 | U-Net | 227.02 | 2.350 | 0.400 | 99.67 | 2.09 |
| T2V-zero [28] | ICCV’23 | U-Net | 209.66 | 2.661 | 0.400 | 20.78 | 2.55 |
| LaVie [71] | Arxiv’23 | U-Net | 166.97 | 2.763 | 0.346 | 77.89 | 2.46 |
| AnimateDiff V3 [22] | ICLR’24 | U-Net | 197.89 | 2.944 | 0.467 | 70.85 | 2.62 |
| VideoCrafter2 [11] | Arxiv’24 | U-Net | 178.45 | 2.753 | 0.433 | 80.10 | 2.68 |
| MCM-MSLION [84] | Arxiv’24 | U-Net | 202.08 | 2.33 | 0.417 | 62.60 | 3.04 |
| MagicTime [83] | Arxiv’24 | U-Net | 257.56 | 1.916 | 0.478 | 81.82 | 3.13 |
| Latte [47] | Arxiv’24 | DiT | 192.12 | 2.111 | 0.363 | 68.68 | 2.20 |
| OpenSora 1.1 [90] | Github’24 | DiT | 195.43 | 2.678 | 0.444 | 73.98 | 2.52 |
| OpenSora 1.2 [90] | Github’24 | DiT | 166.92 | 2.781 | 0.375 | 51.60 | 2.56 |
| OpenSoraPlan v1.1 [41] | Github’24 | DiT | 188.53 | 2.421 | 0.327 | 68.52 | 2.19 |
| EasyAnimate V3 [77] | Arxiv’24 | DiT | 164.30 | 2.713 | 0.349 | 90.54 | 2.32 |
| CogVideoX-2B [81] | Arxiv’24 | DiT | 159.31 | 3.225 | 0.404 | 43.15 | 2.92 |
| OpenSoraPlan v1.1† | Ours | DiT | 185.72 | 2.753 | 0.341 | 49.85 | 3.03 |
| OpenSoraPlan v1.1‡ | Ours | DiT | 180.11 | 2.864 | 0.346 | 70.12 | 3.05 |